# ArxivBench: Can LLMs Assist Researchers in Conducting Research?

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable effectiveness in completing various tasks such as reasoning, translation, and question answering. However the issue of factual incorrect content in LLM-generated responses remains a persistent challenge. In this study, we evaluate both proprietary and open-source LLMs on their ability to respond with relevant research papers and accurate links to articles hosted on the arXiv platform, based on high level prompts. To facilitate this evaluation, we introduce arXivBench, a benchmark specifically designed to assess LLM performance across eight major subject categories on arXiv and five subfields within computer science, one of the most popular categories among them. Our findings reveal a concerning accuracy of LLM-generated responses depending on the subject, with some subjects experiencing significantly lower accuracy than others. Notably, Claude-3.5-Sonnet exhibits a substantial advantage in generating both relevant and accurate responses. And interestingly, most LLMs achieve a much higher accuracy in the Artificial Intelligence sub-field than other sub-fields. This benchmark provides a standardized tool for evaluating the reliability of LLM-generated scientific responses, promoting more dependable use of LLMs in academic and research environments. Our code is open-sourced at https://github.com/arxivBenchLLM/arXivBench and our dataset is available on huggingface at https://huggingface.co/datasets/arXivBenchLLM/arXivBench.
ArxivBench: Can LLMs Assist Researchers in Conducting Research?
Ning Li, Jingran Zhang, Justin Cui University of California, Los Angeles Los Angeles, CA 90095 {ningli23, zhangjingran, justincui}@ucla.edu
@@ 1 Introduction
As large language models (LLMs) are increasingly employed in academic studies and research (Liang et al., 2024; Si et al., 2024), the accuracy of the content they generate need to be inspected carefully, particularly in rigorous academic settings. One of the major challenges with LLMs is their tendency to generate fabricated content without valid references (Ji et al., 2023; Huang et al., 2023), which can undermine the integrity of the research process.
Although LLMs have revolutionized many aspects of daily life, the lack of a standardized benchmark for evaluating their effectiveness in finding research papers, a crucial factor for advancing LLM techniques, is particularly concerning despite that they have been trained on related corpses such as arXiv (Tilwani et al., 2024; Naveed et al., 2023), a major platform for sharing research papers. In this paper, we introduce arXivBench, the first benchmark specifically designed to assess the accuracy of model-generated content in response to research-related prompts on arXiv. ArXivBench comprises a carefully curated dataset of 6,500 prompts from 13 fields, providing a robust platform for evaluating their performance in academic contexts, along with an automatic prompt generation pipeline allowing for the easy scalability of the dataset and an evaluation suite for final performance evaluation.
@@ 2 Related work
Large language models (LLMs) have demonstrated remarkable proficiency across a wide array of tasks, showcasing their versatility and effectiveness in different domains (Abdin et al., 2024; Kenton and Toutanova, 2019; Brown, 2020; Basyal and Sanghvi, 2023; Vaswani, 2017). In mathematical reasoning, LLMs can solve complex problems, provide step-by-step explanations, and assist in advanced problem-solving(Kaddour et al., 2023; Ahn et al., 2024). In the realm of literature creation, these models are capable of generating coherent and creative content, from poetry to narrative storytelling, often emulating diverse writing styles(Gómez-Rodríguez and Williams, 2023; Wang et al., 2024).
One similar work to ours is the REASONS benchmark (Tilwani et al., 2024), which assesses the ability of LLMs to generate accurate citations, but it operates within a relatively constrained framework. Specifically, it prompts LLMs to retrieve the authors of a specific paper title. In contrast, our arXivBench takes a more open-ended approach, evaluating the models’ ability to generate relevant and accurate information across a wider range of topics, such as providing research papers related to a particular field. Instead of limiting the responses to predefined papers, we aim to assess the models’ capacity to generate response that are applicable to broader subject areas. This approach allows for a comprehensive evaluation of the models’ effectiveness.
Retrieval-Augmented Generation (RAG) technology has garnered significant attention and been extensively studied in recent years due to its exceptional ability to mitigate the generation of factually incorrect content by large language models (LLMs) (Gao et al., 2023b; Ram et al., 2023). By integrating a retrieval mechanism, RAG enhances the accuracy of responses by sourcing relevant information from external knowledge, which the model may not have directly learned during training (Lewis et al., 2020). Since RAG is a post-training augmentation technology that’s not applied to most of the LLM APIs, we specifically evaluate LLM performances without RAG.
@@ 3.1 Prompt generation
Table 1: Accuracy rates on arXivBench of eight major subjects, where higher values indicate models generating more relevant and accurate papers. All results were measured with a temperature setting of $0.$ A dash ("-") indicates that no relevant or correct papers were generated. “EESS” refers to Electrical Engineering and Systems Science.
@@ 3.2 Evaluation method
For each prompt, the targeted model is asked for generating three related papers (including titles and arXiv links). The generated links are then used to extract unique arXiv paper IDs, which are cross-referenced with paper information from the Kaggle mirrored arXiv dataset(arXiv.org submitters, 2024). This dataset is updated monthly, ensuring that the information used for cross-referencing remains current and reflects the latest research papers submitted to arXiv.
A selected LLM is introduced to evaluate the model outputs (Zheng et al., 2023; Gilardi et al., 2023; Zeng et al., 2023). First, the LLM determines whether the model’s response is relevant to the prompt. If the response is deemed irrelevant, the LLM returns “no,” and the second evaluation step is skipped. If the response is valid, the LLM proceeds to determine whether the paper title and the arXiv link refer to the same paper, using additional information associated with the paper ID. The LLM returns “yes” if the title and the content fetched by the link matches, “no” if they do not, and “pass” if the LLM is unable to reach a conclusion. If a paper link cannot be found in the Kaggle arXiv dataset, it is automatically classified as incorrect and bypasses the LLM’s judgment.
@@ 4.1 Experiment setup
Table 2: Accuracy rates on arXivBench of subfields with in computer science, where higher values indicate models generating more relevant and accurate papers. All results were measured with a temperature setting of $0.$ A dash ("-") indicates that no relevant or correct papers were generated.
@@ 4.2 Evaluation results
Lastly, we analyze the performance of models within the same family, observing that model size significantly impacts response accuracy, as shown in Table  1 and Table  2. Figure  2 further illustrates how larger models tend to achieve higher accuracy across both major subjects and specific subfields. For example, within the Llama 3.1 family, the large-sized Llama-3.1-405B-Instruct-Turbo consistently outperforms its smaller counterparts. It achieves the highest accuracy in Math (14.87%) and Physics (24.43%), compared to the medium-sized Llama-3.1-70B-Instruct-Turbo, which achieves 8.40% in Math and 23.31% in Physics. The small-sized Llama-3.1-8B-Instruct-Turbo shows significantly lower accuracy, with only 2.38% in Math and 7.92% in Physics. A similar pattern is observed within the Mistral family. The Mistral-large-latest model leads with 21.20% accuracy in Computer Science and 21.99% in Physics, compared to the smaller Mistral-small-latest model, which only achieves 16.73% in Computer Science and 17.12% in Physics.
@@ 5 Conclusion
In this paper, we introduce the first benchmark for accessing the ability of large language models (LLMs) to generate relevant and accurate research responses on arXiv across both major subject categories and specialized subfields. Through our experiments with 15 models, we find significant variability in performance across different subjects, with models like Claude-3.5 Sonnet demonstrating strong results in both broad and specialized domains. However, the overall accuracy remains relatively low, highlighting the challenges LLMs face. Our findings emphasize the need for further refinement of these models to improve their application in academic and research contexts. We hope that arXivBench will serve as a valuable tool for the research community, encouraging the development of more reliable and accurate LLMs for scientific purposes.
@@ A.1 More related work
The ALCE benchmark, introduced by Gao et al. (2023a), does not delve into the academic context or the practical application of large language models (LLMs) in research environments. This creates a gap in evaluating the real-world feasibility of using LLMs for academic purposes, where accuracy and context are crucial. In response to these shortcomings, arXivBench offers a more targeted assessment of LLM performance, particularly in academic and research settings. It rigorously tests models’ abilities to generate accurate and relevant responses across a variety of subjects, providing a clearer understanding of their limitations in handling complex academic tasks.
While RAG emphasizes improving the factual correctness of generated outputs, other benchmarks such as MT-Bench(Zheng et al., 2023) and Chatbot Arena (Chiang et al., 2024) focus on LLMs’ ability to follow instructions and evaluate responses in conversational AI settings. arXivBench complements these efforts by providing a structured evaluation specifically for academic research, where LLMs are tested on their ability to produce accurate and relevant citations without relying on external retrieval systems. In addition, research on models like BERT (Kenton and Toutanova, 2019) and GPT (Brown, 2020) has also paved the way for benchmarking efforts that focus on scaling and improving LLMs for tasks that require high accuracy, such as research citation generation.
@@ A.2 Challenges in research
arXivBench highlights several challenges that need to be addressed before LLMs can reliably support research activities(Li et al., 2024; Talukdar and Biswas, 2024; Radford, 2018). These include overcoming issues such as hallucinations, where models generate plausible but incorrect information, and the need for better context alignment when producing responses(Xu et al., 2024; Rawte et al., 2023). Moreover, it underscores the importance of ensuring that LLMs can reference up-to-date and valid academic sources, a critical aspect in research that often demands real-time accuracy (Zhang et al., 2023; Thoppilan et al., 2022). By addressing these limitations, arXivBench sets the groundwork for future improvements in LLMs to make them more effective and reliable tools for academic research.
@@ A.3 Generator model selection
As introduced in the previous sections, we selected a large language model (LLM) as our prompt generator, a method that has been widely adopted in numerous prior studies (Cui et al., 2024). As outlined in Section 2, this highly advanced LLM is capable of generating creative, concise, and large-scale prompts, which played a crucial role in facilitating the prompt generation for our benchmark. These qualities ensure that the prompts are both diverse and relevant, improving the overall effectiveness of our evaluation. For the model selection, we opted for GPT-4o-mini, an advanced yet cost-efficient model that is well-suited for large-scale generation tasks. Its efficiency and performance make it an ideal choice for our needs, allowing us to generate a substantial number of prompts without compromising on quality or incurring excessive computational costs.
A.4 Model response processor
The model response processor is introduced between the prompting of LLMs and the retrieval of paper information from the Kaggle arXiv dataset. When prompting the LLMs, we provide a detailed set of output instructions in addition to the prompt from the arXivBench dataset as in Fig  4. The processor uses regular expressions to extract paper titles and associated links from the model’s response. As per the instructions, we request that models format titles and links within double square brackets ([[ ]]). However, the models do not always adhere strictly to these formatting guidelines and may place links in single brackets ([ ]) or parentheses instead. Fortunately, our processor is designed to handle these common deviations, ensuring it can accurately parse and retrieve relevant information despite inconsistent formatting.
A.5 Invalid model response
@@
A.6 Judge model selection
As previously introduced, we utilize a large language model (LLM) as our automated judge, as manual evaluation at this scale would be highly labor-intensive and impractical(Zhang et al., 2024). The Llama-3.1-8B model was selected as our judge due to its strong performance as an open-source model. To further assess its accuracy as a judge, we randomly sampled 800 evaluations conducted by the Llama-3.1-8B model. After manually reviewing these evaluations, we determined that the judge achieved an accuracy rate of 98.13%, with a false positive rate (FPR) of 3.61%. This high level of accuracy demonstrates the model’s reliability in evaluating the responses generated by other models.
A.7 More graph on evaluation results
The preceding pages have revealed our finding on the relation between model size and model accuracy, which has been deeply studies before (Kaplan et al., 2020; Hoffmann et al., 2022). We would like to provide an additional graph supporting our finding in Fig  3.
Figure 4: The demo of prompting a LLM

### v2

# ArXivBench: When You Should Avoid Using ChatGPT for Academic Writing

@@ Abstract
Large language models (LLMs) demonstrate strong capabilities in reasoning and question answering, yet their tendency to generate factually incorrect content remains a critical challenge. This study evaluates proprietary and open-source LLMs on generating relevant research papers with accurate arXiv links. Our evaluation reveals critical academic risks: LLMs frequently generate incorrect arXiv links or references to non-existent papers, fundamentally undermining their ability to properly attribute research contributions to the actual authors. We introduce arXivBench, a benchmark specifically designed to assess LLM performance across eight major subject categories on arXiv and five subfields within computer science, one of the most popular categories among them. Our findings show concerning accuracy variations across subjects, with Claude-3.5-Sonnet exhibiting a substantial advantage in generating both relevant and accurate responses. Notably, most LLMs perform significantly better in Artificial Intelligence than other subfields. This benchmark provides a standardized tool for evaluating LLM reliability in scientific contexts, promoting more dependable academic use in research environments. Our code and dataset are available at https://github.com/liningresearch/arXivBench and https://huggingface.co/datasets/arXivBenchLLM/arXivBench.
@@ 1 Introduction
As large language models (LLMs) are increasingly employed in academic studies and research [20, 25], the accuracy of the content they generate need to be inspected carefully, particularly in rigorous academic settings. One of the major challenges with LLMs is their tendency to generate fabricated content without valid references [14, 13], which can undermine the integrity of the research process.
Although LLMs have revolutionized many aspects of daily life, the lack of a standardized benchmark for evaluating their effectiveness in finding research papers, a crucial factor for advancing LLM techniques, is particularly concerning despite that they have been trained on related corpses such as arXiv [28, 21], a major platform for sharing research papers. In this paper, we introduce arXivBench, the first benchmark specifically designed to assess the accuracy of model-generated content in response to research-related prompts on arXiv. ArXivBench comprises a carefully curated dataset of 6,500 prompts from 13 fields, providing a robust platform for evaluating their performance in academic contexts, along with an automatic prompt generation pipeline allowing for the easy scalability of the dataset and an evaluation suite for final performance evaluation.
@@ 2 Related work
Large language models (LLMs) have demonstrated remarkable proficiency across a wide array of tasks, showcasing their versatility and effectiveness in different domains [1, 17, 5, 4, 29]. In mathematical reasoning, LLMs can solve complex problems, provide step-by-step explanations, and assist in advanced problem-solving[15, 2]. In the realm of literature creation, these models are capable of generating coherent and creative content, from poetry to narrative storytelling, often emulating diverse writing styles[11, 30].
One similar work to ours is the REASONS benchmark [28], which assesses the ability of LLMs to generate accurate citations, but it operates within a relatively constrained framework. Specifically, it prompts LLMs to retrieve the authors of a specific paper title. In contrast, our arXivBench takes a more open-ended approach, evaluating the models’ ability to generate relevant and accurate information across a wider range of topics, such as providing research papers related to a particular field. Instead of limiting the responses to predefined papers, we aim to assess the models’ capacity to generate response that are applicable to broader subject areas. This approach allows for a comprehensive evaluation of the models’ effectiveness.
Retrieval-Augmented Generation (RAG) technology has garnered significant attention and been extensively studied in recent years due to its exceptional ability to mitigate the generation of factually incorrect content by large language models (LLMs) [9, 23]. By integrating a retrieval mechanism, RAG enhances the accuracy of responses by sourcing relevant information from external knowledge, which the model may not have directly learned during training [18]. Since RAG is a post-training augmentation technology that’s not applied to most of the LLM APIs, we specifically evaluate LLM performances without RAG.
@@ 3.1 Prompt generation
Table 1: Accuracy rates on arXivBench of eight major subjects, where higher values indicate models generating more relevant and accurate papers. All results were measured with a temperature setting of $0.$ A dash (”-”) indicates that no relevant or correct papers were generated. “EESS” refers to Electrical Engineering and Systems Science.
3.2 Model response processor
The model response processor is introduced between the prompting of LLMs and the retrieval of paper information from the Kaggle arXiv dataset. When prompting the LLMs, we provide detailed output instructions in addition to the prompt from the arXivBench dataset as in Fig  2. The processor uses regular expressions to extract paper titles and associated links from the model’s response. The instructions request that models format titles and links within double square brackets ([[ ]]). However, models do not always follow these guidelines and may place links in single brackets ([ ]) or parentheses instead. Fortunately, our processor handles these common deviations, ensuring it accurately parses and retrieves information despite inconsistent formatting.
@@ 3.3 Evaluation method
For each prompt, the targeted model is asked for generating three related papers (including titles and arXiv links). The generated links are then used to extract unique arXiv paper IDs, which are cross-referenced with paper information from the Kaggle mirrored arXiv dataset[3]. This dataset is updated monthly, ensuring that the information used for cross-referencing remains current and reflects the latest research papers submitted to arXiv.
A selected LLM is introduced to evaluate the model outputs [35, 10, 32]. First, the LLM determines whether the model’s response is relevant to the prompt. If the response is deemed irrelevant, the LLM returns “no,” and the second evaluation step is skipped. If the response is valid, the LLM proceeds to determine whether the paper title and the arXiv link refer to the same paper, using additional information associated with the paper ID. The LLM returns “yes” if the title and the content fetched by the link matches, “no” if they do not, and “pass” if the LLM is unable to reach a conclusion. If a paper link cannot be found in the Kaggle arXiv dataset, it is automatically classified as incorrect and bypasses the LLM’s judgment.
@@ 4.1 Experiment setup
Table 2: Accuracy rates on arXivBench of subfields with in computer science, where higher values indicate models generating more relevant and accurate papers. All results were measured with a temperature setting of $0.$ A dash (”-”) indicates that no relevant or correct papers were generated.
@@ 4.2 Evaluation results
Lastly, we analyze the performance of models within the same family, observing that model size significantly impacts response accuracy, as shown in Table  1 and Table  2. Figure  3 further illustrates how larger models tend to achieve higher accuracy across both major subjects and specific subfields. For example, within the Llama 3.1 family, the large-sized Llama-3.1-405B-Instruct-Turbo consistently outperforms its smaller counterparts. It achieves the highest accuracy in Math (14.87%) and Physics (24.43%), compared to the medium-sized Llama-3.1-70B-Instruct-Turbo, which achieves 8.40% in Math and 23.31% in Physics. The small-sized Llama-3.1-8B-Instruct-Turbo shows significantly lower accuracy, with only 2.38% in Math and 7.92% in Physics. A similar pattern is observed within the Mistral family. The Mistral-large-latest model leads with 21.20% accuracy in Computer Science and 21.99% in Physics, compared to the smaller Mistral-small-latest model, which only achieves 16.73% in Computer Science and 17.12% in Physics.
@@ 5 Conclusion
In this paper, we introduce arXivBench, the first benchmark for assessing LLM accuracy in generating relevant research papers with accurate arXiv links. Evaluating 15 models reveals significant performance variability across subjects, with Claude-3.5 Sonnet demonstrating strong results across domains. We identify critical limitations for academic use: generation of non-existent papers and incorrect arXiv links that prevent proper attribution of research contributions to the actual authors. These findings suggest current models may have reliability limitations for academic writing applications and careful validation remain essential when considering LLMs for scholarly reference generation. We hope arXivBench encourages development of more reliable LLMs while promoting awareness of these limitations in academic contexts.
@@ A.1 More related work
The ALCE benchmark, introduced by Gao et al. [8], does not delve into the academic context or the practical application of large language models (LLMs) in research environments. This creates a gap in evaluating the real-world feasibility of using LLMs for academic purposes, where accuracy and context are crucial. In response to these shortcomings, arXivBench offers a more targeted assessment of LLM performance, particularly in academic and research settings. It rigorously tests models’ abilities to generate accurate and relevant responses across a variety of subjects, providing a clearer understanding of their limitations in handling complex academic tasks.
While RAG emphasizes improving the factual correctness of generated outputs, other benchmarks such as MT-Bench[35] and Chatbot Arena [6] focus on LLMs’ ability to follow instructions and evaluate responses in conversational AI settings. ArXivBench complements these efforts by providing a structured evaluation specifically for academic research, where LLMs are tested on their ability to produce accurate and relevant citations without relying on external retrieval systems. In addition, research on models like BERT [17] and GPT [5] has also paved the way for benchmarking efforts that focus on scaling and improving LLMs for tasks that require high accuracy, such as research citation generation.
@@ A.2 Challenges in research
ArXivBench highlights several challenges that need to be addressed before LLMs can reliably support research activities[19, 26, 22]. These include overcoming issues such as hallucinations, where models generate plausible but incorrect information, and the need for better context alignment when producing responses[31, 24]. Moreover, it underscores the importance of ensuring that LLMs can reference up-to-date and valid academic sources, a critical aspect in research that often demands real-time accuracy [33, 27]. By addressing these limitations, arXivBench sets the groundwork for future improvements in LLMs to make them more effective and reliable tools for academic research.
@@ A.3 Generator model selection
As introduced in the previous sections, we selected a large language model (LLM) as our prompt generator, a method that has been widely adopted in numerous prior studies [7]. As outlined in Section 2, this highly advanced LLM is capable of generating creative, concise, and large-scale prompts, which played a crucial role in facilitating the prompt generation for our benchmark. These qualities ensure that the prompts are both diverse and relevant, improving the overall effectiveness of our evaluation. For the model selection, we opted for GPT-4o-mini, an advanced yet cost-efficient model that is well-suited for large-scale generation tasks. Its efficiency and performance make it an ideal choice for our needs, allowing us to generate a substantial number of prompts without compromising on quality or incurring excessive computational costs.
A.4 Invalid model response
@@
A.5 Judge model selection
As previously introduced, we utilize a large language model (LLM) as our automated judge, as manual evaluation at this scale would be highly labor-intensive and impractical[34]. The Llama-3.1-8B model was selected as our judge due to its strong performance as an open-source model. To further assess its accuracy as a judge, we randomly sampled 800 evaluations conducted by the Llama-3.1-8B model. After manually reviewing these evaluations, we determined that the judge achieved an accuracy rate of 98.13%, with a false positive rate (FPR) of 3.61%. This high level of accuracy demonstrates the model’s reliability in evaluating the responses generated by other models.
A.6 More graph on evaluation results
The preceding pages have revealed our finding on the relation between model size and model accuracy, which has been deeply studies before [16, 12]. We would like to provide an additional graph supporting our finding in Fig  4.